# LEP3: A High-Luminosity $e^{+}e^{-}$ Higgs & Electroweak Factory in the LHC Tunnel

### v2

@@
LEP3: A High-Luminosity $e^{+}e^{-}$ Higgs & Electroweak
Factory in the LHC Tunnel
A possible back-up to the preferred option (FCC-ee and FCC-hh) for the next accelerator for CERN
Updated on May 27, 2025
C. Anastopoulos1, R. Assmann16, A.H. Ball2, O. Bruning3, O. Buchmueller4, T. Camporesi5,15, P. Collier3, J Dainton6,14, G. Davies4, J.R. Ellis3,7, B. Goddard3, L. Gouskos8, G. Hall4, M. Klute9, M. Koratzinos10, G. Landsberg8, K. Long4, L. Malgeri3, F. Maltoni11,17, F. Moortgat3, C. Mariotti12, S. Myers3, J.A. Osborne3, M. Pierini3, D.R. Tovey1, D. Treille3, T.S. Virdee4, N. Wardle4, M. Zanetti13
Contact person: T. Virdee (t.virdee@imperial.ac.uk)
1 University of Sheffield, UK2RAL, STFC, UK3CERN4Imperial College London, UK5LIP, Portugal6Cockroft Inst., Daresbury Lab. UK 7Kings College London, UK8Brown University, USA9Karlsruhe Institute of Technology, Germany10PSI, Switzerland11Louvain University, Belgium12INFN Torino, Italy13University of Padua and INFN Padova, Italy14 University of Liverpool, UK15 Boston University, USA16 GSI Darmstadt, Germany 17University of Bologna and INFN Bologna, Italy
As stated in the 2019 European Strategy for Particle Physics (ESPP), it is of the utmost importance that the HL-LHC upgrade of the accelerator and the experiments be successfully completed in a timely manner. All necessary efforts should be devoted to achieving this goal.
We also recall two of the principal recommendations of the 2019 ESPP for future accelerator initiatives, namely that
  • An electron-positron Higgs factory is the highest priority for the next collider (Recommendation c).
  • Europe, together with its international partners, should investigate the technical and financial feasibility of a future hadron collider at CERN with a centre-of-mass energy of at least 100 TeV and with an electron-positron Higgs and electroweak factory as a possible first stage (Recommendation e).
A major objective in particle physics is always to operate an accelerator that allows a leap of an order of magnitude in the constituent centre-of-mass (CoM) energy with respect to the previous one.
We support FCC-ee and FCC-hh as the preferred option for CERN’s future, as it addresses both of the above recommendations.
The guidance for the 2025 ESPP requests, in addition to the preferred option, the inclusion of “prioritised alternatives to be pursued if the chosen preferred option turns out not to be feasible or competitive”. Proposed alternatives to the preferred FCC option include linear, muon colliders and LHeC accelerators. In response to this request we propose reusing the existing LHC tunnel for an electron-positron collider, called LEP3, as a back-up alternative if the FCC cannot proceed. LEP3 leverages much of the R&D conducted for FCC-ee, offers high-precision studies of Z, W, and Higgs bosons below the $t‑\bar{t}$ threshold, and offers potential physics performance comparable or superior to other fallback options at a lower cost while supporting continued R&D towards a next-generation energy frontier machine.
LEP3 is not intended to compete with the FCC-ee.
@@ 1 Introduction
We discuss here the option of an e+e- Higgs and electroweak factory in the LEP/LHC tunnel – LEP3 - as a possible backup for FCC if the latter is unfeasible for technical or financial reasons. This was previously proposed for the 2013 ESPP but not pursued further [1]. We follow the principal lines developed in that proposal, though the details benefit from studies for FCC-ee [2] and incorporate many components developed for it. Thus R&D carried out for FCC-ee would be maximally utilised. More details are provided in an accompanying document [3].
@@
The baseline anticipates two experiments. During the running programme, around $\approx 4\times 10^{5}\,e^{+}e^{-}$ ZH events would be recorded over six years at $\sqrt{s}=230\,GeV$. LEP3 would also run at the Z and WW thresholds, recording $\approx 2\times 10^{12}$ Z decays over six years at or around $\sqrt{s}=91.2\,GeV$, and $\approx 4\times 10^{7}$ WW events at or around $\sqrt{s}=160\,GeV$ over four years.
@@
Using the FCC-ee cost methodology, the cost to CERN of LEP3, including costs related to civil engineering, LHC removal, LEP3 installation and two new experiments is estimated to be around 3.9 BCHF for all stages. Some 0.9 BCHF from non-CERN contributions in not included in the above number (see Table 3). In the costs presented here we have included the cost of two new experiments at $\approx$0.5 BCHF each with CERN’s contribution amounting to 10% of 1 BCHF. If the cost to teh community has to be limited, the existing ATLAS and CMS experiments could be re-deployed, suitably modified/upgraded.
We consider that the prospective physics performance of LEP3 is fully competitive with other proposed alternatives, to the preferred FCC option, at a lower cost. However, LEP3 is not competitive with FCC-ee.
@@
In an accompanying longer document [3] we have listed questions about the LEP3 option that would need to be addressed if it is to be taken further.
@@ 2 The Main Stages of LEP3
The broad aim is to conduct precision electroweak physics involving the W and Z bosons as well as precision Higgs boson physics. The order proposed above corresponds to increasing the energy monotonically, but another ordering could also be considered (e.g., six years at ZH, six years at and around the Z-peak, and four years at the WW threshold). Running at increasing energy steps has the advantage of spreading out the required expenditure. The duty cycle assumed is the same as for FCC-ee, i.e., 185 days @ 75% efficiency giving an effective running time of $1.2\times 10^{7}$ seconds per year.
@@ 3 Technical Considerations
Luminosity and CoM Energy for ZH Operation: The maximum CoM energy has been chosen to be 230 $GeV$. Although the total Higgs boson cross section peaks at a CoM energy of 250 $GeV$, the luminosity in a circular collider, at a fixed synchrotron radiation power loss, drops rapidly with beam energy, moving the peak rate of Higgs bosons lower (to 235 $GeV$ in the case of LEP3/FCC-ee).
In this update of our submission  [3], we base our parameters on a study made by K Oide and D. Shatilov in 2017 [4] using a preliminary lattice for LEP3, inspired by the FCC-ee design. The Oide/Shatilov study was carried out for four IPs and at $\sqrt{s}=240\,GeV$. Their estimate of the instantaneous luminosity/IP was $1.1\times 10^{34}cm^{-2}s^{-1}$ at $\sqrt{s}=240\,GeV$ and $52\times 10^{34}cm^{-2}s^{-1}/IP$ around the Z-pole. The LEP3 baseline anticipates two IPs and ZH running at $\sqrt{s}=230\,GeV$, leading to an instantaneous luminosity/IP of $1.5\times 10^{34}cm^{-2}s^{-1}$ for ZH running. This increase, from 1.1 to 1.5×$\times 10^{34}cm^{-2}s^{-1}/IP$, is due to two factors of 1.15, corresponding to two experiments and a lowered $\sqrt{s}$ energy, respectively. The Oide/Shatilov study used very strong normal quadrupoles and sextupoles which would lead to a very high power consumption and would need a wide range of hardware site-modifications. For these reasons we propose the use of nested superconducting quadrupoles/sextupoles. This possibility is an option in the FCC-ee design [2][5]. This would result in not only a significantly lower power consumption, but also an increase in instantaneous luminosity as the magnetic bending radius could be 7 % higher (2958m instead of 2755m), reducing the RF required as well as the increasing the luminosity by 7%. We have re-evaluated the LEP3 parameters in the light of this and the Oide/Shatilov study. The parameters from our study are shown in Table 1. We estimate that the maximum RF voltage to be installed will be around 6 GV and the instantaneous luminosity/IP at $\sqrt{s}=230\,GeV$ to be $1.6\times 10^{34}cm^{-2}s^{-1}$. For the Z running, we have changed parameters to respect two aspects: the spin modulation index, a figure of merit for spin depolarization measurements, and the threshold of x-z instability (given by $Q_{s}/\xi_{x}$ ). For the former, the problem was that the momentum compaction factor is much smaller compared to LEP (and similar to FCC). To compensate for that, we have increased $Q_{s}$ by increasing RF voltage. Then, to keep a reasonable beam-beam vertical tune shift, we increased the vertical emittance and decreased the number of bunches. This in turn helped in increasing the bunch length (which is good for x-z instability, resistive wall power, etc.)
@@
To calculate the number of events for the physics analyses we have used a luminosity value of $1.6\times 10^{34}cm^{-2}s^{-1}/IP$ at $\sqrt{s}=230\,GeV$, $6.2\times 10^{34}cm^{-2}s^{-1}/IP$ at or around $\sqrt{s}=160\,GeV$, $40\times 10^{34}cm^{-2}s^{-1}/IP$ at or around $\sqrt{s}=91.2\,GeV$. Hence, for two experiments, around $4.2\times 10^{5}e^{+}e^{-}\rightarrow ZH$ events would be recorded over six years at $\sqrt{s}=230\,GeV,1.7\times 10^{12}Z$ decays over six years at or around $\sqrt{s}=91.2GeV$, and $3.6\times 10^{7}WW$ events at or around $\sqrt{s}=160\,GeV$ over four years. Whereas the rates of Higgs boson production at linear colliders would be similar to that potentially achievable at LEP3 for a single IP, the luminosities are considerably smaller near the Z peak and the WW threshold. Consequently, LEP3 has a much stronger precision electroweak programme, complementing its precision Higgs physics one.
Energy Calibration: It is important for precision electroweak physics to measure the beam energy as precisely as possible for Z and WW running. The contribution of the beam energy uncertainty on the W mass measurement uncertainty at LEP, compared with that of the Z mass mesurement, was a factor 10 higher (25 $MeV$) as no resonant depolarisation measurements were possible at the WW energies (no polarisation was detected). For LEP3, we can use the FCC-ee approach for Z running, based on the knowledge that LEP routinely achieved adequate polarisation levels at the Z. For WW running where no polarization was detected at LEP, precise simulations will be needed to see if adequate levels of polarization can be achieved at energies close to or at the WW threshold. The benefits available with modern technologies, that were not available for LEP1 and LEP2, need to be investigated in addition. For example, much lower than 5% polarisation levels might be sufficient for resonant depolarisation with today’s technologies. Use of modern diagnostics and electronics will thus likely enable energy calibration at higher energies. Failing this, the aim is to use the resonant depolarisation method at the highest energy possible, and extrapolate to the WW threshold. We also propose to install a large amount of RF voltage for the WW running (3GV compared to an energy loss per turn of 1.5 $GeV$. This needs further study.
Crossing Angle and Implications: A “crab-waist” scheme with a significant crossing angle of around 30 $mrad$ (half-angle of 15 $mrad$) is necessary to attain high instantaneous luminosity [6]. Such crossing angles necessitate special optics for the paths of the beams in the vicinity of the IPs. A scheme similar to that deployed for CEPC [7] can be found for LEP3, in which the IP can be kept at the coordinates of the current IPs in ATLAS and CMS, and limiting the spatial excursion of the beams in the straight sections. We have initiated a study with experts to realise this possibility.
@@
Injector: Arguments for a dedicated linac injector, rather than using the SPS as injector, are given in the FCC MTR [2]. Therefore, here we pursue only the linac option. The injection energy is affected by non-uniformities in the bending magnets at the lowest magnetic field i.e. at injection energy. Although the injection energy potentially can be lowered by the ratio of the circumferences, i.e., by a factor of 3.4, we assume a smaller factor, and a higher injection energy of 10 $GeV$. This is to be compared with 20 $GeV$ for FCC-ee. The lower energy allows the linac length to be halved compared to what would be required for FCC-ee, while keeping a similar energy swing as the FCC. We adopt the siting study carried out for FCC-ee to locate the linac injector on the Prevessin site, but the shorter length allows for a more optimal placement minimizing transfer tunnel length. The layout can still be optimized to reduce its length or use existing tunnels.
@@
Table 2: Parameters for the RF system for Z, WW and ZH running
[…]
Civil Engineering Works: The LEP/LHC tunnel was constructed in the 1980s with an assumed lifetime of some 40 years. To prolong its usability, either for LEP3 or as an injector for the future FCC-hh programme, maintenance will be needed in some sections considered to be fragile. It is estimated that this repair corresponds to a section of a length of 0.6 $km$ in Sector 3-4. Furthermore, the straight sections around the two collision IPs will need widening to accommodate the large crossing angle. The length of the tunnel assumed to be widened corresponds to two times 2$\times$270 $m$, with a maximum diameter of  7 $m$. Allowance has been made for some additional cores from the klystron galleries into the tunnel. It is assumed that the existing LHC shafts can be used to carry out the LEP3 civil engineering works. The civil engineering works are estimated to cost around 165 MCHF.
Experiments: Although having two new experiments in the existing ATLAS and CMS caverns may be preferable for many reasons, including a better-adapted physics performance, also use could be made of the two existing LHC general-purpose experiments i.e. ATLAS and CMS suitably modified or upgraded, especially their inner tracking systems. Which option to adopt will require an optimization and considerations concerning the cost, the physics performance and sociology. If this Project proceeds it is probable that a call for experiments would be made and new and/or suitably upgraded existing experiments inevitably would be proposed. A previous study explored the suitability of the CMS experiment for such e+e- operation [8]. The R&D carried out by the ALICE collaboration for its new bendable pixel detector could form a possible basis for LEP3 pixel and tracking detectors. Special luminometers will be required. If existing detectors are used, finer lateral granularity, e.g., for hadronic calorimeters may have to be introduced. The TDRs for the experiments (new or re-deployed) would have to be submitted by the mid-2030s, several years before the start of the shutdown for the transition to LEP3, enabling a head start for construction. The trackers would be designed to be light-weight, with a momentum resolution for 40 $GeV$ charged particles at the $\sim 0.1$% level, and an angular resolution on 40 $GeV$ muons of $\sim 0.1\,mrad$. If ATLAS is reused, modifications to the endcap toroid magnets and/or liquid argon calorimeter may be required to accommodate the final focus of the collider ring. Some level of particle identification can be obtained from the precision timing detectors being installed for HL-LHC running. However, unless the possibility of two new detectors is adopted, probably there is no room to introduce significant $\pi$/K separation capability. We estimate that two new experiments would cost the community some 2$\times$0.5 BCHF. The cost to be borne by CERN is assumed to be 10%, as in the case of FCC-ee costing, ie. 0.1$\times$2$\times$0.5 BCHF = 100 MCHF. This has been included in the cost estimate presented here. In existing detectors, new inner trackers would cost around 100 MCHF per experiment, and we estimate that another 50 MCHF would be needed for all other modifications. It is assumed that a large fraction of the existing experimental collaborations would remain to exploit data from LEP3. If ATLAS and CMS are re-deployed there would be a considerable advantage in using understood detectors, software and analysis tools, and scientists already experienced in physics analysis of data from these existing experiments. For Z running the inner detector would be immersed in a solenoid field of 2T in both the ATLAS and CMS experiments. In CMS the solenoid magnetic field could be raised for higher $\sqrt{s}$ running, e.g., 3 Tesla during ZH running. In ATLAS it is expected that the toroid magnets for muon momentum measurement will continue to operate, possibly at a reduced field. The rest of the ATLAS and CMS detectors should continue to function well during the duration of LEP3 programme.
[…]
@@
Cost Estimate: The capital cost of LEP3 has been evaluated employing the methodology used to estimate the cost of FCC-ee, with appropriate scaling (e.g., by the numbers of components) of the costs of individual FCC-ee items. The error in this estimate would be similar to that for the FCC-ee. Included in this estimate are the costs of the civil engineering works described above, the cost of careful removal and storage of LHC components to allow possible future re-use, and the installation of the LEP3 machine. The latter costs have been estimated by discussions with those involved in the installation of the original LEP accelerator. The total cost to CERN of the LEP3 project, with two new experiments, is estimated to be 3.9 BCHF (see Table 3). The cost is dominated by the cost of the RF system ($\sim 1.5$ BCHF). The next costliest items are the quadrupole/sextupole magnets ($\sim 0.4$ BCHF) and the injector chain ($\sim 0.3$ BCHF). The FCC costing study uses a classification matrix published by the Association for the Advancement of Cost Engineering (AACE).
@@
Table 3: The LEP3 cost to CERN is estimated to be 4785 MCHF using the FCC-ee costing methodology including the additional cost of 900 MCHF to the wordwide particle physics community. The same costs for two existing experiments amount to 4085 MCHF including 270 MCHF respectively. These costs are dominated by the cost of the RF system at over 1.5 BCHF
@@
Schedule:The LEP3 accelerator would be installed after the end of the HL-LHC programme. We estimate that five years will be needed for dismantling the LHC, carrying out civil engineering works and installation and commissioning of the LEP3 accelerator. Assuming the current HL-LHC schedule, LEP3 could be operational in the mid-2040s.
@@
R&D Needs: We believe that around 5% of the capital cost of the Project will have to be spent for the remaining R&D and prototyping before proceeding to construction. This would amount to about 200 MCHF. In addition to the R&D needs for the FCC-ee project outlined in submission  [5], two particular areas of R&D affecting LEP3 are called out: R&D on higher power fundamental power couplers, and on nested superconducting quadrupole/sextupole magnets.
@@
Sustainability: We assume that all building permissions/permits for LEP3 should either be in place or be easy to obtain. Also, the amount of preparatory work and the environmental impact should be minimal. As relatively little construction is required, it is expected that the carbon cost would be a fraction of any new build. Few other significant environmental impacts are anticipated. Nevertheless, benefit would be drawn from any relevant studies in these areas carried out for other proposed projects such as FCC-ee or CLIC. Usage of high-temperature superconducting arc quadrupoles and sextupoles (HTS) magnets would help limit the power consumption compared to using warm magnets.

### v3

@@
Abstract
The 2020 European Strategy for Particle Physics (ESPP) emphasized the critical importance of completing the High-Luminosity LHC (HL-LHC) upgrade of both the accelerator and experiments in a timely manner, identifying it as a top priority for the field. The strategy also established two key recommendations for future accelerator initiatives: (i) the realization of an electron–positron Higgs factory as the highest-priority next collider, and (ii) the investigation, in collaboration with international partners, of the technical and financial feasibility of a hadron collider at CERN with a centre-of-mass energy of at least 100 TeV, potentially preceded by an electron–positron Higgs and electroweak factory. In alignment with these objectives, the Future Circular Collider (FCC) programme—comprising FCC-ee and FCC-hh—represents the preferred path forward for CERN, offering both precision and energy-frontier capabilities. However, the 2025 ESPP update calls for the identification of prioritized alternative options should the preferred FCC pathway prove infeasible or non-competitive. In this context, we propose LEP3, an electron–positron collider reusing the existing LHC tunnel, as a strategic backup to FCC-ee. LEP3 would exploit much of the research and development already carried out for FCC-ee, enabling high-precision studies of the Z, W, and Higgs bosons below the top–antitop production threshold. Combining strong physics potential with reduced cost, LEP3 provides performance comparable or superior to other fallback options—such as linear, muon, or LHeC colliders—while maintaining the technological continuity essential for a future energy-frontier collider. Conceived as a contingency, LEP3 complements, rather than competes with, the FCC-ee proposal.
@@ 1 Introduction
We discuss here the option of an e+e- Higgs and electroweak factory in the LEP/LHC tunnel – LEP3 - as a possible backup for FCC if the latter is unfeasible for technical or financial reasons. This was previously proposed for the 2013 ESPP but not pursued further [1]. We follow the principal lines developed in that proposal, though the details benefit from studies for FCC-ee [2] [3] and incorporate many components developed for it. Thus R&D carried out for FCC-ee would be maximally utilised.
@@
The baseline anticipates two experiments. During the running programme, around $\approx 4\times 10^{5}\,e^{+}e^{-}$ ZH events would be recorded over six years at $\sqrt{s}=230\,GeV$. LEP3 would also run at the Z and WW thresholds, recording $\approx 2\times 10^{12}$ Z decays over six years at or around $\sqrt{s}=91.2\,GeV$, and $\approx 4\times 10^{7}$ WW events at or around $\sqrt{s}=160\,GeV$ over four years. An alternative, shorter, run-plan could comprise 5 years of ZH running, 1 year of WW and 4 years of Z running, with a stop of 2 years after the WW running, especially if higher luminosities are attainable.
@@
Using the FCC-ee cost methodology, the cost to CERN of LEP3, including costs related to civil engineering, LHC removal, LEP3 installation and two new experiments is estimated to be around 3.8 BCHF for all stages. Some 0.7 BCHF from non-CERN contributions is not included in the above number (see Table 3). In the costs presented here we have included the cost of two new experiments at $\approx$ 0.35 BCHF each, as assumed for the FCC-ee experiments, with CERN’s contribution amounting to 10% of 0.7 BCHF. If the cost to the community has to be limited, the existing ATLAS and CMS experiments could be re-deployed, suitably modified/upgraded.
We consider that the prospective physics performance of LEP3 is fully competitive with other proposed alternatives to the preferred FCC option, at a lower cost. However, LEP3 is not competitive with FCC-ee.
@@
Some questions about the LEP3 option would need to be addressed if it is to be taken further.
@@ 2 The Main Stages of LEP3
The broad aim is to conduct precision electroweak physics involving the W and Z bosons as well as precision Higgs boson physics. The order proposed above corresponds to increasing the energy monotonically, but another ordering could also be considered. Running at increasing energy steps has the advantage of spreading out the required expenditure. The duty cycle assumed is the same as for FCC-ee, i.e., 185 days @ 75% efficiency giving an effective running time of $1.2\times 10^{7}$ seconds per year.
@@ 3 Technical Considerations
Instantaneous Luminosity and CoM Energy for ZH Operation: The maximum CoM energy has been chosen to be 230 $GeV$. Although the total Higgs boson cross section peaks at a CoM energy of 250 $GeV$, the luminosity in a circular collider, at a fixed synchrotron radiation power loss, drops rapidly with beam energy, moving the peak rate of Higgs bosons lower (to 235 $GeV$ in the case of LEP3/FCC-ee).
[…]
Further work has been carried out on estimating the instantaneous luminosity using alternative parameters. The estimate given above uses a preliminary lattice developed from the one used for FCC-ee and the extracted numbers are suitably rescaled. The luminosity depends linearly on the maximum achievable vertical beam-beam parameter, which for the case of FCC-ee is 0.134. For LEP3 we have assumed 0.100. The total tune shift in FCC-ee for the case of four interaction points is 0.536, whereas for LEP3 with 2 interaction points it is only 0.200. We can therefore increase the LEP3 beam-beam parameter per IP to values at least as high as FCC-ee and most probably a factor of $\sqrt{2}$ higher, in which case the LEP3 luminosity at 115 GeV per beam will become $3.7\times 10^{34}cm^{-2}s^{-1}$. We estimate that the uncertainty in the instantaneous luminosity at each operating energy is roughly a factor of $\pm 2$.
Another area of improvement, which could considerably decrease the cost of the project, is reducing the beam pipe diameter. The beam pipe diameter is chosen so that beam instabilities are avoided. These instabilities are proportional to the resistive wall impedance of the beam pipe, which grows linearly with the circumference of the collider and inversely with the third root of the bean pipe diameter. Therefore, to keep the same impedance as FCC-ee we can safely reduce the beam pipe diameter by a factor 1.5, going down to 40 $mm$. The cost of the magnets goes with the square of the beam pipe diameter, so the cost saving will be substantial.
We estimate that the maximum RF voltage to be installed will be around 6 GV and the instantaneous luminosity/IP at $\sqrt{s}=230\,GeV$ to be $1.6\times 10^{34}cm^{-2}s^{-1}$. For the Z running, we have changed parameters to respect two aspects: the spin modulation index, a figure of merit for spin depolarization measurements, and the threshold of x-z instability (given by $Q_{s}/\xi_{x}$ ). For the former, the problem was that the momentum compaction factor is much smaller compared to LEP (and similar to FCC). To compensate for that, we have increased $Q_{s}$ by increasing RF voltage. Then, to keep a reasonable beam-beam vertical tune shift, we increased the vertical emittance and decreased the number of bunches. This in turn helped in increasing the bunch length (which is good for x-z instability, resistive wall power, etc.)
@@
To calculate the number of events for the physics analyses we have used a luminosity value of $1.6\times 10^{34}cm^{-2}s^{-1}/IP$ at $\sqrt{s}=230\,GeV$, $6.2\times 10^{34}cm^{-2}s^{-1}/IP$ at or around $\sqrt{s}=160\,GeV$, $40\times 10^{34}cm^{-2}s^{-1}/IP$ at or around $\sqrt{s}=91.2\,GeV$.
Hence, for two experiments, around $4.2\times 10^{5}e^{+}e^{-}\rightarrow ZH$ events would be recorded over six years at $\sqrt{s}=230\,GeV,1.7\times 10^{12}Z$ decays over six years (with the first year of running at 50% efficiency) at or around $\sqrt{s}=91.2GeV$, and $3.6\times 10^{7}WW$ events at or around $\sqrt{s}=160\,GeV$ over four years. Whereas the rates of Higgs boson production at linear colliders would be similar to that potentially achievable at LEP3 for a single IP, the luminosities are considerably smaller near the Z peak and the WW threshold. Consequently, LEP3 has a much stronger precision electroweak programme, complementing its precision Higgs physics one.
Energy Calibration: It is important for precision electroweak physics to measure the beam energy as precisely as possible for Z and WW running. The contribution of the beam energy uncertainty on the W mass measurement uncertainty at LEP, compared with that of the Z mass mesurement, was a factor 10 higher (25 $MeV$) as no resonant depolarisation measurements were possible at the WW energies (no polarisation was detected). For LEP3, we can use the FCC-ee approach for Z running, based on the knowledge that LEP routinely achieved adequate polarisation levels at the Z. For WW running where no polarization was detected at LEP, precise simulations will be needed to see if adequate levels of polarization can be achieved at energies close to or at the WW threshold. The benefits available with modern technologies, that were not available for LEP1 and LEP2, need to be investigated in addition. For example, much lower than 5% polarisation levels might be sufficient for resonant depolarisation with today’s technologies. Use of modern diagnostics and electronics will thus likely enable energy calibration at higher energies. Failing this, the aim is to use the resonant depolarisation method at the highest energy possible, and extrapolate from there to the WW threshold. We also propose to install a large amount of RF voltage for the WW running (3GV compared to an energy loss per turn of 1.5 $GeV$. This area needs further study.
Crossing Angle and Implications: A “crab-waist” scheme with a significant crossing angle of around 30 mradmrad (half-angle of 15 mradmrad) is necessary to attain high instantaneous luminosity [5]. Such crossing angles necessitate special optics for the paths of the beams in the vicinity of the IPs. A scheme similar to that deployed for CEPC [6] can be found for LEP3, in which the IP can be kept at the coordinates of the current IPs in ATLAS and CMS, and limiting the spatial excursion of the beams in the straight sections. We have initiated a study with experts to realise this possibility.
@@
Injector: Arguments for a dedicated linac injector, rather than using the SPS as injector, are given in the FCC MTR [2]. Therefore, here we pursue only the linac option, retained for the FCC-ee. The injection energy is affected by non-uniformities in the bending magnets at the lowest magnetic field i.e. at injection energy. Although the injection energy potentially can be lowered by the ratio of the circumferences, i.e., by a factor of 3.4, we assume a smaller factor, and an injection energy of 10 $GeV$. This is to be compared with 20 $GeV$ for FCC-ee. The lower energy allows the linac length to be halved compared to what would be required for FCC-ee, while keeping a similar energy swing as the FCC. We adopt the siting study carried out for FCC-ee to locate the linac injector on the Prevessin site, but the shorter length allows for a more optimal placement minimizing transfer tunnel length. The layout can still be optimized to reduce the length of transfer tunnels or use existing tunnels.
Synchrotron Power Loss/Turn and RF Requirements: The energy loss/turn for LEP3 operating at 230 $GeV$, with a magnetic bending radius of 2958m, is 5.4 GeVGeV/turn which with a margin leads to a requirement of 6 $GV$ for RF to be installed, compared with 2.1 $GV$ for FCC-ee at $\sqrt{s}=240\,GeV$. Since the RF cost represents the largest item, the optimal choice of RF cavities and their powering is under intense study by CERN’s RF group. Several criteria have to be met simultaneously e.g. power and voltage ratings for cryomodules (CRMs), ratings for power couplers, lateral and longitudinal integration in the LSS etc. The choices presented here have been inspired by the designs retained for the FCC-ee. We believe that the RF system can be further optimised with a view to lowering the overall cost. For WW and ZH running the CERN-RF Group proposes the use of 800 $\mathit{MHz}$, 4-cell RF cavities, similar in design to the FCC-ee 6-cell cavities. These 4-cell cavities should be shorter in length, and the cryomodules shorter by about 1.5m. Due to the high current during Z running the RF Group proposes the use of 800 $\mathit{MHz}$ single cell cavities. Further, they recommend the use of 800 MHz 6-cell cavities for the booster. Costs of installation, removal and replacement of CRMs are included in the overall cost estimate. A summary of the RF cryomodules/cavities is shown in Table 2.
Table 2: Parameters for the RF system for Z, WW and ZH running
A maximum power of 250 kW is assumed for each power coupler. All four even-numbered Long Straight Sections (LSS) will be used for the RF; two for the collider ring (with separators/recombinators) and two for the booster. The LSS have the same length of 545 $m$ with a diameter of 4.4 $m$. The same cavities are used for the electrons and positrons for ZH operation. Adopting the new 4-cell design for RF cryomodules, with a length of 8.45 $m$ for each cryomodule, leads to 52 cryomodules in each of two LSS. Placing one quadrupole with a length of 3.6 $m$ for every 8 cryomodules leads to $52\times 8.45+25m=465m$. Reserving $2\times 40\,m$ for electrostatic separators leads to a total length of 545mm.
Civil Engineering Works: The LEP/LHC tunnel was constructed in the 1980s with an assumed lifetime of some 40 years. To prolong its usability, either for LEP3 or as an injector for the future FCC-hh programme, or for reasons of safety, maintenance will be needed in some sections considered to be fragile. It is estimated that this repair corresponds to a section of a length of 0.6 $km$ in Sector 3-4. Furthermore, the straight sections around the two collision IPs will need widening to accommodate the large crossing angle. The length of the tunnel assumed to be widened corresponds to two times $\times2$70 mm, with a maximum diameter of  7 $m$. Allowance has been made for some additional cores from the klystron galleries into the tunnel. It is assumed that the existing LHC shafts can be used to carry out the LEP3 civil engineering works. The civil engineering works are estimated to cost around 165 MCHF.
Experiments: Although having two new experiments in the existing ATLAS and CMS caverns may be preferable for many reasons, including a better-adapted physics performance, also use could be made of the two existing LHC general-purpose experiments i.e. ATLAS and CMS suitably modified or upgraded, including their inner tracking systems. Which option to adopt will require an optimization and considerations concerning the cost, the physics performance and sociology. If this Project proceeds it is probable that a call for experiments would be made and new and/or suitably upgraded existing experiments inevitably would be proposed. A previous study explored the suitability of the CMS experiment for such e+e- operation [7]. The R&D carried out by the ALICE collaboration for its new bendable pixel detector could form a possible basis for LEP3 pixel and tracking detectors. Special luminometers will be required. If existing detectors are used, finer lateral granularity, e.g., for hadronic calorimeters may have to be introduced. The TDRs for the experiments (new or re-deployed) would have to be submitted by the mid-2030s, several years before the start of the shutdown for the transition to LEP3, enabling a head start for construction. The trackers would be designed to be light-weight, with a momentum resolution for 40 $GeV$ charged particles at the $\sim 0.1$% level, and an angular resolution on 40 $GeV$ muons of $\sim 0.1\,mrad$. If ATLAS is reused, modifications to the endcap toroid magnets and/or liquid argon calorimeter may be required to accommodate the final focus of the collider ring. Some level of particle identification can be obtained from the precision timing detectors being installed for HL-LHC running. However, unless the possibility of two new detectors is adopted, probably there is no room to introduce significant $\pi$/K separation capability.
@@
Machine-Detector Interface: There would be superconducting final quadrupoles near the interaction regions that follow the design outlined in the FCC MTR [2]. Starting at a distance of 2mm from the interaction point, a shielding solenoid surrounds the quadrupole, and a compensating solenoid is located close to the collision point. A luminometer is integrated. A stay-clear cone is defined, above 100 mradmrad, from the IP along the z axis. This corresponds to a line with pseudo rapidity of $|\eta|\sim 3$. Thus the current forward or very forward calorimeters will not be required.
@@
Cost Estimate: The capital cost of LEP3 has been evaluated employing the methodology used to estimate the cost of FCC-ee, with appropriate scaling (e.g., by the numbers of components) of the costs of individual FCC-ee items. The error in this estimate would be similar to that for the FCC-ee. Included in this estimate are the costs of the civil engineering works described above, the cost of careful removal and storage of LHC components to allow possible future re-use, and the installation of the LEP3 machine. The latter costs have been estimated by discussions with those involved in the installation of the original LEP accelerator. The total cost to CERN of the LEP3 project, with two new experiments, is estimated to be 3.8 BCHF (see Table 3). The cost is dominated by the cost of the RF system ($\sim 1.5$ BCHF). The next costliest items are the quadrupole/sextupole magnets ($\sim 0.4$ BCHF) and the injector chain ($\sim 0.3$ BCHF). The FCC costing study uses a classification matrix published by the Association for the Advancement of Cost Engineering (AACE).
@@
Table 3: The LEP3 cost to CERN is estimated to be 3840 MCHF using the FCC-ee costing methodology excluding the additional cost of 700 MCHF to the worldwide particle physics community. The same costs for two existing experiments amount to 3800 MCHF excluding 270 MCHF respectively. These costs are dominated by the cost of the RF system at over 1.5 BCHF
@@
Schedule:The LEP3 accelerator would be installed after the end of the HL-LHC programme. We estimate that five years will be needed for dismantling the LHC, carrying out civil engineering works and installation and commissioning of the LEP3 accelerator. This estimate has been made by some of those previously responsible for the installation, commissioning and removal of the LEP accelerator; and the installation and commissioning of the LHC accelerator. Assuming the current HL-LHC schedule, LEP3 could be operational in the second half of 2040s.
@@
R&D Needs: It is estimated that around 25 MCHF and 40 FTEy, over and above the resources foreseen for the FCC, will be required to advance the project sufficiently to reach a point of decision, anticipated in around 3 years time. The FTEy and financial resources mentioned above are foreseen for prototyping (e.g. of nested superconducting quadrupole and sextupole magnets, fundamental power couplers, etc.) or further studies (including those for RF integration in the existing tunnel, developing a low-emittance LEP3 lattice and optics, booster bypass, injector and its layout, crossing angle, radio-protection related, synchrotron radiation absorption etc.) or consolidation (e.g. of the schedule for the shutdown activities, costing etc.). We believe that around 5% of the capital cost of the Project will have to be spent for the remaining R&D and prototyping before proceeding to construction. This would amount to about 200 MCHF. In addition to the R&D needs for the FCC-ee project outlined in submission  [3], two particular areas of R&D affecting LEP3 are called out above: R&D on higher power fundamental power couplers, and on nested superconducting quadrupole/sextupole magnets.
@@
Sustainability: We assume that all building permissions/permits for LEP3 should either be in place or be easy to obtain. Also, the amount of preparatory work and the environmental impact should be minimal. As relatively little construction is required, it is expected that the equivalent carbon cost would be a fraction of any new build. Few other significant environmental impacts are anticipated. Nevertheless, benefit would be drawn from any relevant studies in these areas carried out for other proposed projects such as FCC-ee or CLIC. Usage of high-temperature superconducting arc quadrupoles and sextupoles (HTS) magnets would help limit the power consumption compared to using warm magnets.